OpenStudy (anonymous):

Solve each system of equations 3x-2y=2 10x-3y=-8

OpenStudy (stamp):

Would you prefer to solve this via substitution or elimination?

OpenStudy (anonymous):

okay 1st you would sove for y in the 1st one

OpenStudy (anonymous):

anyway to get the answer

OpenStudy (anonymous):

so that would be -3x+2=2y

OpenStudy (anonymous):

then simplify to -3/2x+1=y, then you would substitute -3/2x+1 for y in the second problem

OpenStudy (anonymous):

so 10x-3(-3/2x+1)=-8

OpenStudy (anonymous):

do you get that so far?

OpenStudy (anonymous):

wait just a sec

OpenStudy (anonymous):

alrightie(:

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

sorry got distracted with other things